Harper Jones is evaluating an investment project will have an installed cost of $518,297. The cash flows over the 4-year life of the investment are projected to be $287,636, $203,496, $63,024 and $132,191, respectively. What is the NPV of this project if the discount rate is 6.5 percent?

Group of answer choices

$37,306

$55,318

$86,127

$128,415

$169,193
Answer #1

Year Net Cash flow Present value factor Present Value In A B = (1/1.065)^n C = A*B 0 -5,18,297.00 1 -5,18,297.00 1 2,87,636.0